How many mL of 12 M NH3 should be used to make 750. mL of 0.400 M NH3
What volume of 0.125 M HCl can be made from 25.0 mL of 12 M HCl?
A 10.00 g sample of NaOH is to be used to make a NaOH solution having a molarity of 0.100 M. What is the total volume of this solution in mL that can be made from this sample?
1
Expert's answer
2020-03-05T07:56:37-0500

1) In 750 mL of 0.4 M NH3 there are 0.4⋅750/1000 = 0.3 moles of NH3

(1/12)⋅0.3=0.025 L = 25 mL

2) There are 12⋅25.0/1000 = 0.3 moles of HCl in 25 mL

0.3/0.125 = 2.4 L

3) n(NaOH) = 10/40 = 0.25 moles

0.25/0.1 = 2.5 L

ANSWER: 25 mL; 2.4 L; 2.5 L.